OpenStudy (anonymous):

Dave and Jane Wells have a new rectangular driveway. The perimeter of the driveway is 168 feet. the length is 12 feet longer than three times the width. What are the dimensions of the driveway?

OpenStudy (anonymous):

P=L+W L= 12+3W (12+3W)+ W =168 12+4W=168 W=168-12/4 W=39 L=12+3(39) L=129 length = 129 width = 39 Check: P= 129+39=168ft